how to solve: 4^(3x-1)=3^(x-1)
i can't rember how to solve for x.
i am lost.

Answer by edjones(8007)

4^(3x-1)=3^(x-1)
log[4](4^3x-1)=log[3](3^(x-1))/log[3](4) Change of base formula from 3 to 4 on the right side.
3x-1=(x-1)/1.2619
3.7856x-1.2619=x-1
2.7856x=.2619
x=.094...
.
Ed